A metal sphere of radius a has charge QA, it is enclosed by a conducting spherical shell of radius b that has charge QB. Determine:
a) the potential VA  of the inner sphere
b) the potential VB of the outer sphere
c) potential difference between the two sphere (VA-VB)
d) under what conditions i) VA=0 ii)VB =0 iii) VA=VB

Grade:12

2 Answers

Vikas TU
6 years ago
As we have the charge QA and the sweep a 
As we ol realize that V = kQ/r 
At that point capability of internal circle = kQA/a 
Potential contrast of external circle (VB) = kQB/b 
Dhananjay verma
6 years ago
Potential of sphere A is - self potential and potential due to be
Va= KQa/a +KQb/b
Also the potential of sphere B can be find out.
(1) Va=0=Vb , when charge on b and a is zero.
(2) Va=VB , when charge is only on sphere b.